# Free Descriptions of Convex Sets

### v2

@@ 6.2 Fitting Freely-Described Convex Functions to Data
We now present an algorithm for the free convex regression problem from Section 1.1.4. Recall that in this problem we are given data $\{(x_{i},\phi_{i})\in\mathbb{V}_{n_{i}}\oplus\mathbb{R}\}$ in finitely-many dimensions $n_{i}$ corresponding to a sequence of vector spaces $\{\mathbb{V}_{n}\}$, and our objective is to identify a sequence of convex functions $\{f_{n}:\mathbb{V}_{n}\rightarrow\mathbb{R}\}$ such that $f_{n_{i}}(x_{i})\approx\phi_{i}$ in the dimensions in which data is available. We tackle this problem by endowing $\{\mathbb{V}_{n}\}$ with the structure of a consistent sequence of $\{\mathsf{G}_{n}\}$-representations, and choosing description spaces $\{\mathbb{W}_{n}\},\{\mathbb{U}_{n}\}$ that yield a finitely-parametrized family of freely-described and compatible convex functions which we fit to the given data. We explain how to perform this fitting below in two key steps.
@@
$f_{n}(x;\alpha,\beta,\lambda)$ $=\inf_{\begin{subarray}{c}t\geq 0\\ y\in\mathbb{W}_{n}\end{subarray}}t+\lambda\|y\|\quad\textrm{ s.t. }\sum_{i=1}^{d_{A}}\alpha_{i}A_{n}^{(i)}x+\sum_{j=1}^{d_{B}}\beta_{j}B_{n}^{(j)}y+tu_{n}\in\mathcal{K}_{n}$ (P)
$=\sup_{z\in\mathbb{U}_{n}}-\sum_{i=1}^{d_{A}}\alpha_{i}\langle z,A_{n}^{(i)}x\rangle\quad\textrm{ s.t. }\left\|\sum_{j=1}^{d_{B}}\beta_{j}(B_{n}^{(j)})^{*}z\right\|\leq\lambda,\ \langle z,u_{n}\rangle\leq 1,\ z\in\mathcal{K}_{n}^{*},$ (D)
[…]
@@
$\min_{\begin{subarray}{c}\varepsilon\in\mathbb{R}^{D}_{+}\\ \alpha\in\mathbb{R}^{d_{A}},\beta\in\mathbb{R}^{d_{B}},\lambda\geq\lambda_{\min}\\ \{(t_{i},y_{i})\},\{z_{i}\}\end{subarray}}$ $\|\varepsilon\|_{\ell_{2}}\qquad\mathrm{s.t.}$ (Regress)
@@
The quantity $n_{0}$ in Algorithm 1 is governed by the choice of description spaces and leads to a tradeoff between the richness of the parametric family and the dimensions of the available data. The data we are given only contains information about $\{f_{n}\}_{n\leq\max_{i}n_{i}}$ which, in turn, only depends on $\{\{A_{n}^{(i)}\}_{n\leq\max_{i}n_{i}}\}_{i=1}^{d_{A}}$ and $\{\{B_{n}^{(j)}\}_{n\leq\max_{i}n_{i}}\}_{j=1}^{d_{B}}$. If $n_{0}$ in Algorithm 1 is strictly larger than the maximum dimension $\max_{i}n_{i}$ in which data is available, then the number of distinct basis elements for $n\leq\max_{i}n_{i}$ might be strictly smaller than $d_{A}$ and $d_{B}$, respectively, in which case our parameters $\alpha,\beta$ are not identifiable from such low-dimensional data. Imposing compatibility decreases the bound on $n_{0}$, thus facilitating free convex regression from lower-dimensional data. More broadly, even when $\max_{i}n_{i}\geq n_{0}$, it is of interest to investigate the landscape of (Regress).
@@ 6.3 Numerical Results
The first function we approximate is the $\ell_{\pi}$ norm $\|x\|_{\pi}=\left(\sum_{i}|x_{i}|^{\pi}\right)^{1/\pi}$, which is not SDP-representable because $\pi$ is irrational. We view the $\ell_{\pi}$ norm as defined on the sequence $\mathscr{V}=\{\mathbb{R}^{n}\}$ with $\mathsf{G}_{n}=\mathsf{B}_{n}$ from Example 1.5. It satisfies both intersection and projection compatibility. We choose description spaces $\mathscr{W}=\mathscr{U}=\{\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 1}\mathbb{R}^{2n+1})=\mathbb{S}^{2n+2}\}$ as in Example 6.2(c), with the corresponding PSD cones $\{\mathcal{K}_{n}=\mathbb{S}^{2n+2}_{+}\}$. We used 50 data points in $\mathbb{R}^{n}$ for $n_{i}\in\{1,2\}$. When we impose compatibility on our family of functions, we use $n_{0}=2=\max_{i}n_{i}$ in Algorithm 1. If we do not impose compatibility and search over all freely-described (possibly incompatible) sequences, we take $n_{0}=4$ which is the presentation degree of $\mathscr{W}\otimes\mathscr{U}$, and which strictly exceeds $\max_{i}n_{i}$. The constraints in the fitting program (Regress) do not depend on some of the entries of $\alpha,\beta$ in this case, and we set such entries to zero. This highlights the advantage of imposing compatibility mentioned in Remark 6.3–it allows us to uniquely identify a free description from lower-dimensional data.
@@ 7 Conclusions and Future work
We developed a systematic framework to study convex sets that can be instantiated in different dimensions using representation stability, as well as a computational method to parametrize such sets and fit them to data. We did so by formally defining free descriptions of convex sets and compatibility conditions relating sets in different dimensions. We then proved a number of structural results pertaining to free descriptions, namely, characterizing such descriptions certifying compatibility; giving conditions on fixed-dimensional descriptions ensuring their extendability to free ones; and studying infinite-dimensional limits of freely-described sequences of sets. We further used representation stability to systematically derive constant-sized descriptions for sequences of invariant sections of PSD and relative entropy cones. Finally, we developed an algorithm to computationally parametrize and search over free descriptions to fit them to data. Our work can be viewed as identifying and exploiting a new point of contact between representation stability and convex geometry through conic descriptions of convex sets.
@@
  1. fnum@enumiitem(Computational algebra)
@@
  3. fnum@enumiitem(Constructing descriptions)
    Given a sequence of convex sets instantiable in any relevant dimension, can we systematically construct freely-described, and possibly compatible, approximations for the sequence? When are approximations derived from sums-of-squares hierarchies such as [38] free and certify compatibility?
  4. fnum@enumiitem(Complexity)
    Is there a systematic framework to study the smallest possible size of a free description for a given sequence of sets, extending the slack operator-based approach for fixed convex sets [87]?
  5. fnum@enumiitem(Free separation)
@@
  6. fnum@enumiitem(Statistical inference)
@@ Acknowledgements
    We thank Mateo Díaz for helpful discussions and suggestions for improving our implementation, Jan Draisma for suggesting relevant references, and Jorge Garza Vargas for helpful discussions regarding approximately finite-dimensional algebras. The authors were supported in part by Air Force Office of Scientific Research grants FA9550-22-1-0225 and FA9550-23-1-0070, and by National Science Foundation grant DMS 2113724.
@@ Appendix A Representations of Categories
    A (real) representation of a category $\mathcal{C}$, also called a $\mathcal{C}$-module, is a functor $\mathcal{C}\to\mathrm{Vect}_{\mathbb{R}}$ from $\mathcal{C}$ to the category of real vector spaces.
@@
      The category corresponding to Example 1.5 with $\mathsf{G}_{n}=\mathsf{B}_{n}$ (resp., $\mathsf{D}_{n}$) is $\mathcal{C}=\mathrm{FI}_{BC}$ (resp., $\mathcal{C}=\mathrm{FI}|_{D}$) defined in [50], whose objects are the sets [±$[\pm n]\coloneqq\{\pm 1,\ldots,\pm n\}$ for $n\in\mathbb{N}$ and whose morphisms are injections $f\colon[\pm n]\hookrightarrow[\pm N]$ satisfying $f(-i)=-f(i)$ (and reverse evenly-many signs if $\mathsf{G}_{n}=\mathsf{D}_{n}$).
@@
      The category corresponding to the graphon sequence in Section 4.7 is the opposite category $\mathcal{C}=\mathcal{P}_{2}^{\mathrm{op}}$ of the category $\mathcal{P}_{2}$ with objects $[2^{n}]$ and morphisms which are $2^{N-n}$-to-one surjections $[2^{N}]\to[2^{n}]$, or equivalently, partitions of $[2^{N}]$ into $2^{n}$ equal parts.
@@
    1. missingfnum@enumiiitem(Noetherianity)
@@
    2. missingfnum@enumiiitem(Tensor products)
@@
    We say that a category $\mathcal{C}$ satisfies property (TFG) if Tensor products of free $\mathcal{C}$-modules are Free and satisfy the Generation degree bound in Theorem A.5.
@@
    Define the consistent sequence $\mathbb{Z}_{n}=\mathbb{R}[\mathsf{G}_{n}]\left(A_{d_{W}}^{\dagger}\mathbb{W}_{d_{W}}\right)\subseteq\mathbb{V}_{n}$ if $n\geq d_{W}$ and $\mathbb{Z}_{n}=0$ otherwise, where $A_{d_{W}}^{\dagger}$ is the pseudoinverse of $A_{d_{W}}$. Note that $\{\mathbb{Z}_{n}\}$ is generated in degree $d_{W}$. Moreover, $A_{n}^{-1}(\mathbb{W}_{n})=\ker A_{n}+\mathbb{Z}_{n}$. Indeed, we have $A_{n}A_{d_{W}}^{\dagger}=A_{d_{W}}A_{d_{W}}^{\dagger}=\mathrm{id}_{\mathbb{U}_{d_{W}}}$ because $\{A_{n}\}$ is a surjective morphism, hence $A_{n}(\ker A_{n}+\mathbb{Z}_{n})=A_{n}(\mathbb{Z}_{n})=\mathbb{R}[\mathsf{G}_{n}]\mathbb{W}_{d_{W}}=\mathbb{W}_{n}$. Conversely, if $A_{n}x\in\mathbb{W}_{n}=\mathbb{R}[\mathsf{G}_{n}]\mathbb{W}_{d_{W}}$ then we can write $A_{n}x=\sum_{i}g_{i}w_{i}$ for $g_{i}\in\mathsf{G}_{n}$ and $w_{i}\in\mathbb{W}_{d_{W}}$. Then $\hat{x}=\sum_{i}g_{i}A_{d_{W}}^{\dagger}w_{i}\in\mathbb{Z}_{n}$ and $A_{n}(x-\hat{x})=0$, so $x\in\ker A_{n}+\mathbb{Z}_{n}$. Since $\ker\mathscr{A}$ is generated in degree $d$ and $\{\mathbb{Z}_{n}\}$ is generated in degree $d_{W}$, their sum is generated in degree $\max\{d,d_{W}\}$. ∎
@@
    Let =$\mathscr{F}=\{\mathbb{F}_{n}\}$ be a free $\mathcal{C}$-module generated in degree $d$ and let =$\mathscr{B}=\{B_{n}\}\colon\mathscr{F}\to\mathscr{V}$ be a surjective morphism whose kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ is generated in degree $k$. The composition $\mathscr{F}\xrightarrow{\mathscr{B}}\mathscr{V}\xrightarrow{\mathscr{A}}\mathrm{im}\mathscr{A}$ is a surjective morphism from the free $\mathcal{C}$-module $\mathscr{F}$ whose kernel is $\mathscr{B}^{-1}(\mathrm{ker}\mathscr{A})$ and is generated in degree $\max\{d,k\}=k$ by Lemmas A.9 and A.10. ∎
@@
    By Proposition A.7, the $\mathcal{C}$-module $\mathscr{V}^{\otimes|\lambda|}$ is generated in degree $d|\lambda|$ and presented in degree $d(|\lambda|-1)+k$. Let $\mathsf{S}_{|\lambda|}$ act on each $\mathbb{V}_{n}^{\otimes|\lambda|}$ by permuting its factors $\sigma\cdot(v_{1}\otimes\cdots\otimes v_{|\lambda|})=v_{\sigma(1)}\otimes\cdots\otimes v_{\sigma(|\lambda|)}$, which is an orthogonal action commuting with the emebddings $\mathbb{V}_{n}^{\otimes|\lambda|}\subseteq\mathbb{V}_{n+1}^{\otimes|\lambda|}$. In this way, any element $c_{\lambda}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ defines a morphism $c_{\lambda}\colon\mathscr{V}^{\otimes|\lambda|}\to\mathscr{V}^{\otimes|\lambda|}$ such that $c_{\lambda}^{*}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ is also a morphism. If $c_{\lambda}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ is the Young symmetrizer corresponding to partition $\lambda$, then $\mathrm{im}\,c_{\lambda}=\mathbb{S}^{\lambda}\mathscr{V}$. The result follows from Proposition A.11. ∎
@@
    1. missingfnum@enumiiitem(Sums)
@@
    2. missingfnum@enumiiitem(Images and kernels)
@@
    1. missingfnum@enumiiitem(Tensors)
@@
    2. missingfnum@enumiiitem(Schur functors)

### v3

@@ 6.2 Fitting Freely-Described Convex Functions to Data
We now present an algorithm for the dimension-free convex regression problem from Section 1.1.4. Recall that in this problem we are given data $\{(x_{i},\phi_{i})\in\mathbb{V}_{n_{i}}\oplus\mathbb{R}\}$ in finitely-many dimensions $n_{i}$ corresponding to a sequence of vector spaces $\{\mathbb{V}_{n}\}$, and our objective is to identify a sequence of convex functions $\{f_{n}:\mathbb{V}_{n}\rightarrow\mathbb{R}\}$ such that $f_{n_{i}}(x_{i})\approx\phi_{i}$ in the dimensions in which data is available. We tackle this problem by endowing $\{\mathbb{V}_{n}\}$ with the structure of a consistent sequence of $\{\mathsf{G}_{n}\}$-representations, and choosing description spaces $\{\mathbb{W}_{n}\},\{\mathbb{U}_{n}\}$ that yield a finitely-parametrized family of freely-described and compatible convex functions which we fit to the given data. We explain how to perform this fitting below in two key steps.
@@
$f_{n}(x;\alpha,\beta,\lambda)$ $=\inf_{\begin{subarray}{c}t\geq 0\\ y\in\mathbb{W}_{n}\end{subarray}}t+\lambda\|y\|\qquad\textrm{ s.t. }\sum_{i=1}^{d_{A}}\alpha_{i}A_{n}^{(i)}x+\sum_{j=1}^{d_{B}}\beta_{j}B_{n}^{(j)}y+tu_{n}\in\mathcal{K}_{n}$ (P)
$=\sup_{z\in\mathbb{U}_{n}}-\sum_{i=1}^{d_{A}}\alpha_{i}\langle z,A_{n}^{(i)}x\rangle\qquad\textrm{ s.t. }\left\|\sum_{j=1}^{d_{B}}\beta_{j}(B_{n}^{(j)})^{*}z\right\|\leq\lambda,\ \langle z,u_{n}\rangle\leq 1,\ z\in\mathcal{K}_{n}^{*},$ (D)
where the primal and dual programs above are equal since Slater’s condition is satisfied by our choice of $u_{n}$. Here $\|\cdot\|$ is the norm on $\mathbb{W}_{n}$ induced by its inner-product and the parameter $\lambda$ is chosen to be positive; the purpose of this term in (P) is to prevent numerical issues arising in our subsequent regression procedure (see (Regress)). Note that this is indeed a finitely-parametrized (by $d_{A}+d_{B}+1$ parameters) infinite sequence of functions. Also note that if we require compatibility in Algorithm 1, then the sequence $\{f_{n}\}$ is intersection (and if desired, projection) compatible for any value of the parameters, since it is the sequence of gauge functions of a correspondingly compatible sequence of sets (see Section 1.3).
@@
$\min_{\begin{subarray}{c}\varepsilon\in\mathbb{R}^{D}_{+}\\ \alpha\in\mathbb{R}^{d_{A}},\beta\in\mathbb{R}^{d_{B}},\lambda\geq\lambda_{\min}\\ \{(t_{i},y_{i})\},\{z_{i}\}\end{subarray}}$ $\|\varepsilon\|_{\ell_{2}}\hskip 18.49988pt\mathrm{s.t.}$ (Regress)
@@
The quantity $n_{0}$ in Algorithm 1 is governed by the choice of description spaces and leads to a tradeoff between the richness of the parametric family and the dimensions of the available data. The data we are given only contains information about $\{f_{n}\}_{n\leq\max_{i}n_{i}}$ which, in turn, only depends on $\{\{A_{n}^{(i)}\}_{n\leq\max_{i}n_{i}}\}_{i=1}^{d_{A}}$ and $\{\{B_{n}^{(j)}\}_{n\leq\max_{i}n_{i}}\}_{j=1}^{d_{B}}$. If $n_{0}$ in Algorithm 1 is strictly larger than the maximum dimension $\max_{i}n_{i}$ in which data is available, then the number of distinct basis elements for $n\leq\max_{i}n_{i}$ might be strictly smaller than $d_{A}$ and $d_{B}$, respectively, in which case our parameters $\alpha,\beta$ are not identifiable from such low-dimensional data. Imposing compatibility decreases the bound on $n_{0}$, thus facilitating dimension-free convex regression from lower-dimensional data. More broadly, even when $\max_{i}n_{i}\geq n_{0}$, it is of interest to investigate the landscape of (Regress).
@@ 6.3 Numerical Results
The first function we approximate is the $\ell_{\pi}$ norm $\|x\|_{\pi}=\left(\sum_{i}|x_{i}|^{\pi}\right)^{1/\pi}$, which is not SDP-representable because $\pi$ is irrational. We view the $\ell_{\pi}$ norm as defined on the sequence $\mathscr{V}=\{\mathbb{R}^{n}\}$ with $\mathsf{G}_{n}=\mathsf{B}_{n}$ from Example 1.5. It satisfies both intersection and projection compatibility. We choose description spaces $\mathscr{W}=\mathscr{U}=\{\mathrm{Sym}^{2}(\mathrm{Sym}^{\leq 1}\mathbb{R}^{2n+1})=\mathbb{S}^{2n+2}\}$ as in Example 6.2(c), with the corresponding PSD cones $\{\mathcal{K}_{n}=\mathbb{S}^{2n+2}_{+}\}$. We used 50 data points in $\mathbb{R}^{n}$ for $n_{i}\in\{1,2\}$. When we impose compatibility on our family of functions, we use $n_{0}=2=\max_{i}n_{i}$ in Algorithm 1. If we do not impose compatibility and search over all freely-described (possibly incompatible) sequences, we take $n_{0}=4$ which is the presentation degree of $\mathscr{W}\otimes\mathscr{U}$, and which strictly exceeds $\max_{i}n_{i}$. The constraints in the fitting program (Regress) do not depend on some of the entries of $\alpha,\beta$ in this case, and we set such entries to zero. This highlights the advantage of imposing compatibility mentioned in Remark 6.3–it allows us to uniquely identify a dimension-free description from lower-dimensional data.
@@ 7 Conclusions and Future Work
We developed a systematic framework to study convex sets that can be instantiated in different dimensions using representation stability, as well as a computational method to parametrize such sets and fit them to data. We did so by formally defining dimension-free descriptions of convex sets and compatibility conditions relating sets in different dimensions. We then proved a number of structural results pertaining to dimension-free descriptions, namely, characterizing such descriptions certifying compatibility; giving conditions on fixed-dimensional descriptions ensuring their extendability to dimension-free ones; and studying infinite-dimensional limits of freely-described sequences of sets. We further used representation stability to systematically derive constant-sized descriptions for sequences of invariant sections of PSD and relative entropy cones. Finally, we developed an algorithm to computationally parametrize and search over dimension-free descriptions to fit them to data. Our work can be viewed as identifying and exploiting a new point of contact between representation stability and convex geometry through conic descriptions of convex sets.
@@
  1. item(Computational algebra)
@@
  3. item(Constructing descriptions)
    Given a sequence of convex sets instantiable in any relevant dimension, can we systematically construct freely-described, and possibly compatible, approximations for the sequence? When are approximations derived from sums-of-squares hierarchies such as [38] dimension-free and certify compatibility?
  4. item(Complexity)
    Is there a systematic framework to study the smallest possible size of a dimension-free description for a given sequence of sets, extending the slack operator-based approach for fixed convex sets [86]?
  5. item(Dimension-free separation)
@@
  6. item(Statistical inference)
@@ Acknowledgements
    We thank Mateo Díaz for helpful discussions and suggestions for improving our implementation, Jan Draisma for suggesting relevant references, and Jorge Garza Vargas for helpful discussions regarding approximately finite-dimensional algebras. We are grateful to reviewer X for carefully reading our paper and offering numerous detailed suggestions for its improvement. The authors were supported in part by Air Force Office of Scientific Research grants FA9550-22-1-0225 and FA9550-23-1-0070, and by National Science Foundation grant DMS 2113724.
@@ Appendix A Representations of Categories
    A (real) representation of a category $\mathcal{C}$, also called a $\mathcal{C}$-module, is a functor $\mathcal{C}\to\mathrm{Inn}_{\mathbb{R}}$ from $\mathcal{C}$ to the category whose objects are real inner product spaces and whose morphisms are isometries. [^7] Representations of categories are typically defined as functors to the category of all vector spaces over a given field. The results in this section continue to hold for functors to the category of all real vector spaces if we allow consistent sequences consisting of non-orthogonal representations and non-isometric linear maps between them.
@@
    If $n\leq N\leq M$, we compose morphisms by setting
    $(g_{M}\varphi_{M-1}\cdots\varphi_{N})\circ(g_{N}\varphi_{N-1}\cdots\varphi_{n})=(g_{M}g_{N})\varphi_{M-1}\cdots\varphi_{n},$
    which is well-defined and satisfies the necessary axioms since $\varphi_{N},\ldots,\varphi_{M-1}$ are $\mathsf{G}_{N}$-equivariant when acting on the vector spaces in $\mathscr{V}$.
@@
      The category corresponding to Example 1.5 with $\mathsf{G}_{n}=\mathsf{B}_{n}$ (resp., $\mathsf{D}_{n}$) is $\mathcal{C}=\mathrm{FI}_{BC}$ (resp., $\mathcal{C}=\mathrm{FI}|_{D}$) defined in [50], whose objects are the sets $\pm n]\coloneqq\{\pm 1,\ldots,\pm n\}$ for $n\in\mathbb{N}$ and whose morphisms are injections $f\colon[\pm n]\hookrightarrow[\pm N]$ satisfying f(i)=f(i)f(-i)=-f(i) (and reverse evenly-many signs if $\mathsf{G}_{n}=\mathsf{D}_{n}$).
@@
      The category corresponding to the graphon sequence in Section 4.7 is the opposite category $\mathcal{C}=\mathcal{P}_{2}^{\mathrm{op}}$ of the category $\mathcal{P}_{2}$ with objects $2^{n}]$ and morphisms which are $2^{N-n}$-to-one surjections $2^{N}]\to[2^{n}]$, or equivalently, partitions of $2^{N}]$ into $2^{n}$ equal parts.
@@
    1. fnum@enumiiitem(Noetherianity)
@@
    2. fnum@enumiiitem(Tensor products)
@@
    We say that a category $\mathcal{C}$ satisfies property (TFG) if:
    1. (a)
      tensor products of free $\mathcal{C}$-modules are free;
    2. (b)
@@
    Suppose $\{y_{i}\}\subseteq\mathbb{W}_{d_{W}}$ are generators for $\mathscr{W}$. Since $A_{d_{W}}$ is surjective, we can find $\{x_{i}\}\subseteq\mathbb{V}_{d_{W}}$ such that $A_{d_{W}}x_{i}=y_{i}$ for all $i$. Also let $\{z_{j}\}\subseteq\ker A_{d}$ be generators for $\ker\mathscr{A}$. We claim that $\{x_{i}\}\cup\{z_{j}\}\subseteq A_{\max\{d,d_{W}\}}^{-1}(\mathbb{W}_{\max\{d,d_{W}\}})$ generate $\mathscr{A}^{-1}(\mathscr{W})$. Indeed, if $n\geq\max\{d,d_{W}\}$ and $x\in A_{n}^{-1}(\mathbb{W}_{n})$, then we can write $A_{n}x=\sum_{i}\alpha_{i}g_{i}y_{i}$ for some $\alpha_{i}\in\mathbb{R}$ and $g_{i}\in\mathsf{G}_{n}$, in which case $x-\sum_{i}\alpha_{i}g_{i}x_{i}\in\ker A_{n}$ and thus, we have $x=\sum_{i}\alpha_{i}g_{i}x_{i}+\sum_{j}\beta_{j}g_{j}z_{j}$ for some $\beta_{j}\in\mathbb{R}$ and $g_{j}\in\mathsf{G}_{n}$. ∎
@@
    Let $\mathscr{F}=\{\mathbb{F}_{n}\}$ be a free $\mathcal{C}$-module generated in degree $d$ and let $\mathscr{B}=\{B_{n}\}\colon\mathscr{F}\to\mathscr{V}$ be a surjective morphism whose kernel $\mathscr{K}=\{\mathbb{K}_{n}\}$ is generated in degree $k$. The composition $\mathscr{F}\xrightarrow{\mathscr{B}}\mathscr{V}\xrightarrow{\mathscr{A}}\mathrm{im}\mathscr{A}$ is a surjective morphism from the free $\mathcal{C}$-module $\mathscr{F}$ whose kernel is $\mathscr{B}^{-1}(\mathrm{ker}\mathscr{A})$ and is generated in degree $\max\{d,k\}=k$ by Lemmas A.9 and A.10. ∎
@@
    By Proposition A.7, the $\mathcal{C}$-module $\mathscr{V}^{\otimes|\lambda|}$ is generated in degree $d|\lambda|$ and presented in degree $d(|\lambda|-1)+k$. Let $\mathsf{S}_{|\lambda|}$ act on each $\mathbb{V}_{n}^{\otimes|\lambda|}$ by permuting its factors $\sigma\cdot(v_{1}\otimes\cdots\otimes v_{|\lambda|})=v_{\sigma(1)}\otimes\cdots\otimes v_{\sigma(|\lambda|)}$, which is an orthogonal action commuting with the embeddings $\mathbb{V}_{n}^{\otimes|\lambda|}\subseteq\mathbb{V}_{n+1}^{\otimes|\lambda|}$. In this way, any element $c_{\lambda}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ defines a morphism $c_{\lambda}\colon\mathscr{V}^{\otimes|\lambda|}\to\mathscr{V}^{\otimes|\lambda|}$ such that $c_{\lambda}^{*}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ is also a morphism. If $c_{\lambda}\in\mathbb{R}[\mathsf{S}_{|\lambda|}]$ is the Young symmetrizer corresponding to partition $\lambda$, then $\mathrm{im}\,c_{\lambda}=\mathbb{S}^{\lambda}\mathscr{V}$. The result follows from Proposition A.11. ∎
@@
    1. fnum@enumiiitem(Sums)
@@
    2. fnum@enumiiitem(Images and kernels)
@@
    1. fnum@enumiiitem(Tensors)
@@
    2. fnum@enumiiitem(Schur functors)